# Unified Approach to Power-Efficiency Trade-Off of Generic Thermal Machines

### v1

@@ Abstract
Due to the diverse functionalities of different thermal machines, their optimization relies on a case-by-case basis, lacking unified results. In this work, we propose a general approach to determine power-efficiency trade-off relation (PETOR) for any thermal machine. For cases where cycle (of duration $\tau$) irreversibility satisfies the typical $1/\tau$-scaling, we provide a unified PETOR which is applicable to heat engines, refrigerators, heat exchangers and heat pumps. It is shown that, some typical PETORs, such as those for low-dissipation Carnot cycles (including heat engine and refrigerator cycles) and the steady-state heat engines operating between finite-sized reservoirs are naturally recovered.
@@
Introduction.— Finite-time thermodynamics bridges the gap between the idealized quasi-static thermodynamic cycle described by equilibrium thermodynamics and the practical thermal machines operating within finite cycle times [1, 2]. Quantitatively, the performance of finite-time thermodynamic cycles is completely characterized by power-efficiency trade-off relation (PETOR) [3, 4, 5, 6, 7, 8, 9, 10], a consequence of the energy dissipation inherent in the cycle’s irreversibilities [5, 11, 12, 13]. In previous studies, various PETORs have been derived with different theoretical frameworks, such as PETOR for endo-reversible Carnot heat engine [3], PETOR for linear irreversible heat engines [4, 9], and PETOR for low-dissipation heat engine [6, 7]. Besides, one key parameter of finite-time thermal machines, the efficiency at maximum power [14, 15, 16, 17, 18], stands out as a typical point on the power-efficiency tradeoff curve [7, 19, 20, 21], delineating the energy conversion efficiency when the power is maximized.
Figure 1: The normalized power-efficiency trade-off relation of four typical thermal machines operating between two reservoirs of temperatures $T_{\rm{c}}/T_{\rm{h}}=0.5$. The blue dashed (red dash-dotted) curve represents Eq. (1) with the dissipation ratio $\delta\rightarrow 1$($\delta\rightarrow 0$), the normalized reversible efficiency $\tilde{\eta}=1$ is plotted with the green dotted line, and the shadow area marks the available operation region of the machine.
Regarding the quest for completeness and universality in the description and optimization for practical thermal machines, there are two critical limitations in the existing studies: i) In contrast to the rich studies on PETORs for heat engines, discussions on the PETORs for refrigerators and heat pumps remain relatively scarce [22, 23]; ii) Some studies attempt to develop a generalized approach for optimizing various thermal machines [24, 25, 26, 27], although the proposed methods are general in some sense, specific optimizations still require the identification the particular machine before implementation. Note that parameters characterizing thermal machine’s performance, such as output power and energy conversion efficiency, are determined by specific design functions, with entirely different definitions for different machines. For instance, concerning energy conversion efficiency, the coefficient of performance(COP) for refrigerators [28] and heat pumps [29] the efficiency for heat engines have distinct definitions. In this sense, it seems unlikely to obtain a unified PETOR for general thermodynamic cycles, necessitating optimize different thermal machines on a case-by-case basis.
Our purpose in this Letter is to formulate a unified power-efficiency trade-off applicable to any thermal machine, free from restriction imposed by inconsistent definitions of performance parameters for different machines, thus overcoming the aforementioned limitations. By connecting the geometric lower bound of irreversibility for non-equilibrium thermodynamic cycles [11] with thermal machine performance parameters, which are uniformly described with a pair of independent sign functions associated with the cycle direction and machine function. We succinctly obtain the PETOR for a generic thermal machine (whose specific features are determined a coefficient $\chi$) operating in the slow-driving regime as
where $\tilde{P}\equiv P/P_{\rm{max}}$ and $\tilde{\eta}\equiv\eta/\eta_{\rm{rev}}$ are the normalized power and efficiency, respectively. $P_{\rm{max}}$ is the maximum output power of the machine whose reversible efficiency is denoted as $\eta_{\rm{rev}}$. This result is applicable to heat engines, refrigerators, heat pump, and heat exchanger cycles operating between arbitrary number of heat reservoirs, thereby contributing a unifying perspective on power-efficiency trade-off in finite-time thermodynamics. In Fig. 1, we illustrated the PETORs in different regions of $\theta$ and $\gamma$ (see Tab. 1).
Table 1: Cycle quantities of typical thermal machines
@@
Heat Engine $+1$ $+1$ $W$ $Q^{\rm{h}}$ $W/Q^{\rm{h}}$ $W/\tau$
Heat Exchanger $+1$ $-1$ $Q^{\rm{c}}$ $Q^{\rm{h}}$ $Q^{\rm{c}}/Q^{\rm{h}}$ $Q^{\rm{c}}/\tau$
Refrigerator $-1$ $-1$ $Q^{\rm{c}}$ $W$ $Q^{\rm{c}}/W$ $Q^{\rm{c}}/\tau$
Heat Pump $-1$ $+1$ $Q^{\rm{h}}$ $W$ $Q^{\rm{h}}/W$ $Q^{\rm{h}}/\tau$
Unified thermodynamic performance tradeoff.— Typical thermal machines, including heat engines, heat exchangers [30], refrigerators, and heat pumps, are designed to convert energy in various forms according to actual demands. Among them, heat engines harness the temperature difference between hot and cold heat reservoirs to perform work. Both refrigerators and heat pumps create heat flow from cold reservoirs to hot reservoirs through work input, one for cooling and the other for heating. In general sense, the energy conversion efficiency of any thermal machine can be defined as $\eta\equiv\rm{O}/\rm{I}$, where $\rm{O}$ and $\rm{I}$ are respectively the energy exchange at the input end and output end of the machine, as illustrated in Tab. 1 [30]. It should be mentioned here that "output" is categorized from the perspective of the machine’s function, not necessarily the direction of energy flow. For instance, for a general heat engine cycle illustrated in Fig. 2(a), its output function is performing work externally, $\rm{O}$ represents work output $W$ and $\rm{I}$ represents total heat absorption $Q^{\rm{h}}=\sum_{j}Q^{{\rm{h}},j}$ with $Q^{{\rm{h}},j}$ the heat absorption from the $j$-th hot reservoir (We do not restrict the number $N$ of reservoirs, and the following discussion is applicable to any situation with $N\geq 2$). In this case, the direction of energy flow at the output side of the machine is also from the machine to the external environment. However, for a refrigerator cycle, where the output function is absorbing heat from the low-temperature end, namely the direction of energy flow is from the surroundings to the refrigerator. For thermal machines operating within finite cycle time, by dividing $\rm{O}$($\rm{I}$) into reversible part $\rm{O}_{\rm{rev}}$($\rm{I}_{\rm{rev}}$) and irreversible part $A_{\rm{o}}$($A_{\rm{i}}$), we further express the machine efficiency as
@@
Here, $\rm{O}_{\rm{rev}}$, $\rm{I}_{\rm{rev}}$, and $A_{\rm{o}/\rm{i}}$ are all positive. $\theta=\pm 1$ is a sign function determined by the direction of the cycle, while $\gamma=\pm 1$ is another sign function determined by the specific function of the machine with a given $\theta$, as illustrated in Tab. 1. In Fig 2(b), a clockwise ($\theta>0$) and a counterclockwise ($\theta<0$) finite-time Carnot cycle between two reservoirs of temperatures $T_{\rm{h}}>T_{\rm{c}}$ is depicted, where $J$-axis indicates the direction of energy flow exiting the cycles. In such a geometric representation, the greater the output power of the cycle, the further away it is from the plane of $J=0$.
Figure 2: (a) schematic of a heat engine operates among multiple heat reservoirs which is represented by red and blue circles. The engine extracts from supplied reservoirs and releases to others. (b) Temperature-entropy ($T-S$) diagram of clockwise and counterclockwise finite-time Carnot cycles. The orientation of the $J$-axis represents the direction of energy flow from the cycles to the heat reservoirs. (c) EMPs of heat engine ($\tilde{\eta}_{\rm{MP}}\geq 0.5$) and refrigerator ($\tilde{\eta}_{\rm{MP}}\leq 0.5$) as functions of $\delta$. In this example with two reservoirs of temperatures $T_{\rm{c}}<T_{\rm{h}}$, dash-dotted curve, dashed curve, and solid curve are plotted with $T_{\rm{c}}/T_{\rm{h}}=0.1,0.5,0.8$, respectively.
Table 2: Dissipation and its ratio associated with input and output ends of typical thermal machines
$A_{\rm{o}}$ $A_{\rm{i}}$ $A$ $A_{\rm{o}}/A$ $A_{\rm{i}}/A$
Heat Engine $W_{\rm{ir}}$ $Q_{\rm{ir}}^{\rm{h}}$ $W_{\rm{ir}}$ $1$ $\delta$
@@
Refrigerator $Q_{\rm{ir}}^{\rm{c}}$ $W_{\rm{ir}}$ $W_{\rm{ir}}$ $1-\delta$ $1$
Heat Pump $Q_{\rm{ir}}^{\rm{h}}$ $W_{\rm{ir}}$ $W_{\rm{ir}}$ $\delta$ $1$
The dissipation $W_{\rm{ir}}$ in the work output/input process, denoted as $A$, is contributed by all heat transfer processes together [31]. For typical thermal machines, we illustrate the dissipation in different processes in Tab. 2, where the dissipation ratios of different processes are determined with the definition $\delta\equiv Q_{\rm{ir}}^{\rm{h}}/W_{\rm{ir}}\in[0,1]$. Taking use of the cycle parameters, we can unify $A_{\rm{o}}$ and $A_{\rm{i}}$ in terms of $A$ as
@@
Then Eq. (2) becomes
@@
$A=\frac{\chi\left(\rm{O}_{\rm{rev}}-\rm{I}_{\rm{rev}}\eta\right)}{\theta\gamma\left(1-\chi\eta\right)\delta^{\frac{1+\theta}{2}}},$ (5)
@@
Physically, it is reasonable to assume that $A$ has a cycle duration $\tau$-dependent lower bound, i.e., $A\geq A_{\rm{min}}(\tau)$. By further noticing the output power of the machine is $P=\rm{O}/\tau$, Eq. (5) yields
$\frac{\chi\left(\rm{O}_{\rm{rev}}-\rm{I}_{\rm{rev}}\eta\right)}{\theta\gamma\left(1-\chi\eta\right)\delta^{\frac{1+\theta}{2}}}\geq A_{\rm{min}}(\tau)=A_{\rm{min}}\left(\frac{\rm{O}}{P}\right).$ (7)
@@
Eliminating ${\rm{O}}=\rm{O}_{\rm{rev}}-\theta\gamma A_{\rm{o}}$ from the right side of the above equation with Eqs. (3) and (5), an general inequality with respect to $P$ and $\eta$ is obtained.
Specifically, in the slow-driving regime, the minimal energetic dissipation is proved to be $A_{\rm{min}}=\mathcal{L}^{2}/\tau$ [11], where $\mathcal{L}$ is the thermodynamic length in thermodynamic geometry theory. This low bound for irreversibility is achieved when the driving protocol of the cycle maintains uniform speed along the thermodynamic length trajectory (in the parametric space defined by thermodynamic state variables) [11, 32, 33]. With this typical $1/\tau$-scaling, Eq. (7) becomes
@@
$\frac{\chi\left(\rm{O}_{\rm{rev}}-\rm{I}_{\rm{rev}}\eta\right)}{\theta\gamma\left(1-\chi\eta\right)\delta^{\frac{1+\theta}{2}}}\geq\frac{\mathcal{L}^{2}P}{\rm{O}}=\frac{\mathcal{L}^{2}P}{\rm{O}_{\rm{rev}}-\theta\gamma A_{\rm{o}}},$ (8)
substituting Eqs. (3) and (5) into which, by straightforward calculation, we find the Unified trade-off relation between power and efficiency as

### v2

# Unified approach to power-efficiency trade-off relations of generic thermal machines

@@ Abstract
We present a general framework for determining the power-efficiency trade-off relations across arbitrary thermal machines, addressing the lack of unified optimization results stemming from their diverse functionalities (e.g., heat engines, refrigerators, and heat pumps). For time-dependent cycle irreversibility $A(\tau)$ following a $\tau^{-\alpha}$ power law, where $\alpha$ is an interaction-dependent parameter, we show that engineering the interactions between thermal machines and reservoirs enables control over the trade-off relations, with the efficiency at maximum power approaching Carnot efficiency as $\alpha$ increases. Setting $\alpha=1$ naturally recovers typical low-dissipation regime results. Additionally, we derive the first power-efficiency trade-off for finite-time quantum adiabatic Otto machines with $\tau^{-2}$-scaling. This work establishes a unified constraint for thermodynamic cycles across non-equilibrium regimes, facilitating consistent optimization of diverse thermal devices in practice.
I Introduction
@@
Finite-time thermodynamics bridges the gap between idealized quasi-static cycles in equilibrium thermodynamics and real-world thermal machines operating within finite cycle times, standing as a widely employed non-equilibrium thermodynamic theoretical approach of substantial practical importance [1, 2, 3]. A central challenge in this field is characterizing the performance of thermal machines quantitatively through the power-efficiency trade-off relation (hereinafter referred to as the trade-off relation for brevity) [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], which arises from energy dissipation due to irreversibilities [7, 14, 15, 16]. Previous studies have derived trade-off relations for specific models, such as the endo-reversible Carnot heat engine [5, 13], linear irreversible heat engines [6, 11], and low-dissipation heat engines [8, 9]. Beyond these conventional trade-off relations, the thermodynamic trade-offs in diverse practical scenarios, including information-energy conversion [17, 12], particle separation [18, 19], battery charging [20], and biochemical motors [21], have recently garnered growing research interest. Besides, one key parameter of finite-time thermal machines, efficiency at maximum power (EMP) [4, 22, 23, 24, 25, 26], stands out as a typical point on the trade-off relation [9, 27, 28, 29], delineating the energy conversion efficiency when the power is maximized.
However, research on the trade-off relation in thermal machines still suffers from two critical limitations. The first lies in the lack of universality across thermal machine types. Existing trade-off relations are predominately derived for heat engines, while investigations into refrigerators and heat pumps remain notably insufficient [30, 31]. More critically, no universal trade-off relation has been established to unify different thermal machine categories. The core obstacle stems from the absence of a cohesive framework capable of describing these diverse machines—rooted in the inconsistency of their performance metrics [32, 33, 34, 35]. For instance, energy conversion efficiency, a key parameter for characterizing trade-offs, is defined distinctly across devices: heat engines rely on "efficiency", while refrigerators and heat pumps use the "coefficient of performance (COP)" [36, 37]. This divergence in metrics makes a one-size-fits-all trade-off relation for general thermodynamic cycles seemingly unachievable, forcing optimization to be conducted on a case-by-case basis for different thermal machines. The second limitation concerns the restricted applicable regime of the trade-off relations. Due to the equivalence among three core thermal machine models—endoreversible, linear irreversible, and low-dissipation [38, 39], nearly all existing trade-off relations [5, 6, 8, 9, 11, 16, 13] are tied to the typical $1/\tau$-irreversibility ($\tau$ denotes cycle duration) [40, 9, 15]. Nevertheless, recent studies have revealed that the intrinsic specificity of working substances [41, 42, 43] or the diversity of system-reservoir interactions [44] can modify the low-dissipation characteristics of heat engines, thereby significantly affecting their finite-time performance. Consequently, developing a universal trade-off relation that transcends the $1/\tau$-irreversibility regime is urgently needed to optimize practical thermal devices in broader, real-world operating regimes.
In this work, we tackle the aforementioned challenges by deriving a unified trade-off relation: universal for all thermal machine (any function) and beyond the low-dissipation regime. By connecting the $\tau$-dependent irreversibility $A(\tau)\propto\tau^{-\alpha}$ [45, 44] to the performance parameters of thermal machines, we succinctly derive the trade-off relation for a generic thermal machine. Our general result naturally recovers the previous methods when the dimensionless parameter $\alpha$, which depends on the system–reservoir interaction [44], is specified, as will be shown later. The concrete contents of this paper are organized as follows: In Sec. II, we develop a mathematical framework for the general thermal machine, where the sign functions are introduced and the generalized efficiency and dissipation are defined. In Sec. III.1, we derive the trade-off relation between power and efficiency from the unified description of thermal machines, together with the lower bound of dissipation. In Sec. III.2, we further investigate the EMP of the thermal machines. Two typical non-equilibrium operating regime of finite-time thermal machines are discussed in detail in Sec. IV to demonstrate the validity of our general results. Finally, we draw our conclusions in Sec. V.
II Unified description of thermal machines
In this section we employ the sign function as a powerful tool to characterize the efficiency of the thermal machines. By decomposing the system energy into reversible and irreversible contributions, we derive a unified relation between efficiency and dissipation.
Table 1: Cycle quantities of thermal machines
@@ II Unified description of thermal machines
Heat Engine +1+1 +1+1 $W$ $Q^{\rm{h}}$ $W/Q^{\rm{h}}$ $W/\tau$
Heat Exchanger +1+1 1-1 $Q^{\rm{c}}$ $Q^{\rm{h}}$ $Q^{\rm{c}}/Q^{\rm{h}}$ $Q^{\rm{c}}/\tau$
Refrigerator 1-1 1-1 $Q^{\rm{c}}$ $W$ $Q^{\rm{c}}/W$ $Q^{\rm{c}}/\tau$
Heat Pump 1-1 +1+1 $Q^{\rm{h}}$ $W$ $Q^{\rm{h}}/W$ $Q^{\rm{h}}/\tau$
Thermal machines are designed to convert energy in various forms according to actual demands, including heat engines, heat pumps, refrigerators, and heat exchangers. Although heat exchangers are typically applied to individual heat transfer processes rather than full thermodynamic cycles in engineering practice, we include them in Table 1 for completeness. Among them, heat engines harness the temperature difference between hot and cold heat reservoirs to perform work. Both refrigerators and heat pumps create heat flow from cold reservoirs to hot reservoirs through work input, one for cooling and the other for heating. In general sense, the energy conversion efficiency of any thermal machine can be broadly cognitively defined as $\eta\equiv\rm{O}/\rm{I}$, where $\rm{O}$ and $\rm{I}$ are respectively the energy exchange at the input side and output side of the machine, as illustrated in Tab. 1. It should be mentioned here that "output" is categorized from the perspective of the machine’s function, not necessarily the direction of energy flow. For instance, for a general heat engine cycle illustrated in Fig. 1(a), its output function is performing work externally, $\rm{O}$ represents work output $W$ and $\rm{I}$ represents total heat absorption $Q^{\rm{h}}=\sum_{j}Q^{{\rm{h}},j}$ with $Q^{{\rm{h}},j}$ the heat absorption from the $j$-th hot reservoir (We do not restrict the number $N$ of reservoirs, and the following discussion is applicable to any situation with $N\geq 2$[46]. In this case, the direction of energy flow at the output side of the machine is also from the machine to the external environment. However, for a refrigerator cycle, where the output function is absorbing heat from the low-temperature side, namely the direction of energy flow is from the surroundings to the refrigerator.
For thermal machines operating within finite cycle time, by dividing $\rm{O}$($\rm{I}$) into its scalar-valued reversible part $\rm{O}_{\rm{rev}}$($\rm{I}_{\rm{rev}}$) and irreversible part $A_{\rm{o}}$($A_{\rm{i}}$) on the basis of the second law of thermodynamics, we further express the machine efficiency as
@@
Here, $\rm{O}_{\rm{rev}}$, $\rm{I}_{\rm{rev}}$, and $A_{\rm{o}/\rm{i}}$ are all positive. The efficiency expression presented above stems from the following physical intuition: $\theta=\pm 1$ is a sign function determined by the direction of the cycle, while $\gamma=\pm 1$ is another sign function determined by the specific function of the machine output with a given $\theta$, as illustrated in Tab. 1.
Specifically, after determining the cycle direction $\theta$, the direction of heat flow is also determined, either from the high- to the low-temperature heat reservoir or vice versa. For such a three-side machine (comprising two heat sides and one work side), we may define the side where heat flow converges as the net sink. If the net sink aligns with the machine’s desired output side, the condition $\theta\gamma<0$ holds. This also leads to monotonic behavior in the optimization curve of the heat pump. For example, in heat exchanger and heat pump where $\theta\gamma<0$ , the output cold reservoir(or hot reservoir for heat pump) is just defined at the net sink of the energy flow path; in contrast, for heat engine and refrigerator, the output corresponds to the intermediate tap terminal of the energy flow, respectively. Quantitatively, in Fig. 1(b), a clockwise ($\theta>0$) and a counterclockwise ($\theta<0$) finite-time Carnot cycle between two reservoirs of temperatures $T_{\rm{h}}>T_{\rm{c}}$ is depicted, where $J$-axis indicates the direction of energy flow exiting the cycles. In such a geometric representation, the greater the output power of the cycle, the further away it is from the plane of $J=0$.
Figure 1: (a) schematic of a heat engine operates among multiple heat reservoirs which is represented by red and blue circles. The engine extracts from supplied reservoirs and releases to others. (b) Temperature-entropy (TST-S) diagram of clockwise and counterclockwise finite-time Carnot cycles. The orientation of the $J$-axis represents the direction of energy flow from the cycles to the heat reservoirs.
Table 2: Dissipation and its ratio associated with input and output sides of thermal machines
$A_{\rm{o}}$ $_{\rm{i}} AA$ $A_{\rm{o}}/A$ $A_{\rm{i}}/A$
Heat Engine $W_{\rm{ir}}$ $Q_{\rm{ir}}^{\rm{h}}$ $W_{\rm{ir}}$ 11 $\delta$
@@
Refrigerator $Q_{\rm{ir}}^{\rm{c}}$ $W_{\rm{ir}}$ $W_{\rm{ir}}$ $\delta 11$
Heat Pump $Q_{\rm{ir}}^{\rm{h}}$ $W_{\rm{ir}}$ $W_{\rm{ir}}$ $\delta$ 11
The dissipation $W_{\rm{ir}}$ in the work output/input process, denoted as $A$, is contributed by all heat transfer processes together. This additivity of dissipation is jointly induced by the first and second laws of thermodynamics. Taking a heat engine cycle as an example, we have $Q^{\rm{h}}-Q^{\rm{c}}=W$ and $Q^{\rm{h}}_{\rm{rev}}-Q^{\rm{c}}_{\rm{rev}}=W_{\rm{rev}}$. Subtract the above two equations yields $\left(Q^{\rm{h}}_{\rm{rev}}-Q^{\rm{h}}\right)+\left(Q^{\rm{c}}-Q^{\rm{c}}_{\rm{rev}}\right)=W_{\rm{rev}}-W$, which can be further simplified as $Q^{\rm{h}}_{\rm{ir}}+Q^{\rm{c}}_{\rm{ir}}=W_{\rm{ir}}$. For thermal machines, we illustrate the dissipation in different processes in Tab. 2, where the dissipation ratios of different processes are determined with the definition $\delta\equiv Q_{\rm{ir}}^{\rm{h}}/W_{\rm{ir}}\in[0,1]$. In actual implementations, $\delta$ can be tuned via the thermal properties of system-reservoir interaction strength. For an ideal gas, we derive $\delta$ as a function of system parameters in Appendix A via thermodynamic geometry approach. Taking use of the cycle parameters, we can unify output dissipation $A_{\rm{o}}$ and input dissipation $A_{\rm{i}}$ in terms of $A$ as
@@
III Universal thermodynamic bounds
In this section, within the theoretical framework developed above for thermal machines of different functionalities and under the assumption of a lower bound on dissipation, we derive the trade-off relations and the EMPs. We then analyze the influence of the parameter $\alpha$ on this relation and illustrate the results with examples of a heat engine and a refrigerator. Notably, for sufficiently large $\alpha$ there exists a parameter regime in which operation at maximum power can coincide with efficiency approaching the reversible limit.
III.1 Trade-off relations between power and efficiency
Substituting Eq. (2) into Eq. (1) yields
@@ III.1 Trade-off relations between power and efficiency
$A=\frac{\chi\left(\rm{O}_{\rm{rev}}-\rm{I}_{\rm{rev}}\eta\right)}{\theta\left(1-\chi\eta\right)\delta^{\frac{1+\theta}{2}}},$ (4)
@@
Physically, it is reasonable to assume that $A$ has a cycle duration $\tau$-dependent lower bound [47, 48, 49, 50, 51, 15, 52, 53], which has been explored in theoretical and experimental aspects i.e., $A\geq A_{\rm{min}}(\tau)$. By further noticing that the output power of the machine is $P=\rm{O}/\tau$, Eq. (4) yields
$\frac{\chi\left(\rm{O}_{\rm{rev}}-\rm{I}_{\rm{rev}}\eta\right)}{\theta\left(1-\chi\eta\right)\delta^{\frac{1+\theta}{2}}}\geq A_{\rm{min}}(\tau)=A_{\rm{min}}\left(\frac{\rm{O}}{P}\right).$ (6)
@@
Eliminating ${\rm{O}}=\rm{O}_{\rm{rev}}-\theta\gamma A_{\rm{o}}$ from the right side of the above equation with Eqs. (2) and (4), a general inequality with respect to $P$ and $\eta$ is obtained.
As a concrete yet general case, we consider $A_{\rm{min}}=\Sigma\tau^{-\alpha}$ with $\Sigma$ denoting a dissipation coefficient [45, 41, 44]. This form captures a wide range of scenarios, including finite-time isothermal processes in the slow-driving regime [25, 15] and finite-time quantum adiabatic processes with energy oscillations neglected [41, 51]. In this case, Eq. (6) becomes
@@
$\frac{\chi\left(\rm{O}_{\rm{rev}}-\rm{I}_{\rm{rev}}\eta\right)}{\theta\left(1-\chi\eta\right)\delta^{\frac{1+\theta}{2}}}\geq\Sigma\left(\frac{P}{\rm{O}}\right)^{\alpha}=\Sigma\frac{P^{\alpha}}{(\rm{O}_{\rm{rev}}-\theta\gamma A_{\rm{o}})^{\alpha}},$ (7)
substituting Eqs. (2) and (4) into which and performing straightforward calculation yields the main result of this work